# MS-Occ: Multi-Stage LiDAR-Camera Fusion for 3D Semantic Occupancy Prediction

### v1

@@ III-A2 Semantic-aware Module
As shown in Fig. 4, given the voxel features $F_{\mathrm{L}}\in\mathbb{R}^{D\times H\times W\times C}$, only a limited number of voxels are actually occupied by LiDAR point clouds. For each occupied voxel at location $(x_{i},y_{i},z_{i})$, where $x_{i}\in\left\{0,1,\cdots,W-1\right\}$, $y_{i}\in\left\{0,1,\cdots,H-1\right\}$, and $z_{i}\in\left\{0,1,\cdots,D-1\right\}$, the voxel center is projected onto the depth-aware image features of each view, denoted as $F_{\mathrm{I}}^{\mathrm{geo}}$. The corresponding 2D image features $F_{\mathrm{I}}^{\mathrm{fet}}$ are then extracted at these projected points and combined with the original occupied voxel features $F_{\mathrm{L}}^{\mathrm{occ}}$ through element-wise addition. The resulting aggregated features $F_{\mathrm{L}}^{\mathrm{add}}$ serve as queries and interact with the multi-scale image features $F_{\mathrm{I}}^{\mathrm{fet}}$ of the respective image views via a deformable attention module [23, 3]. This process is performed independently for each image view:
@@
$\mathtt{DA}\left(u,r,e\right)=\sum_{i=\mathrm{1}}^{N_{\mathrm{head}}}W_{i}\sum_{j=\mathrm{1}}^{N_{\mathrm{key}}}A_{ij}\cdot W_{i}^{\prime}e\left(r+\Delta r_{ij}\right).$ (4)
@@
Here, $u$ and $e$ denote the queries and keys, respectively, $W_{i}$ and $W_{i}^{\prime}$ are learnable projection matrices, $A_{ij}$ represents the attention weights, $r$ is the 2D reference point onto which the voxel center is projected, and $\Delta r_{ij}$ denotes the positional offset. Due to the overlapping fields of multiple views, a voxel may be projected onto several image views. To consolidate the updated features across views, max-pooling is applied, resulting in the output feature $\mathtt{F_{\mathrm{L}}^{\mathrm{sem}}=\left\{F_{\mathrm{L},k}^{\mathrm{sem}}\right\}_{k=\mathrm{1}}^{\mathit{N}_{\mathrm{v}}}}$, where $\mathtt{\mathit{N}_{\mathrm{v}}}$ is the number of occupied voxels. Each updated voxel feature $F_{\mathrm{L},k}^{\mathrm{sem}}$ is defined as:
$F_{\mathrm{L},k}^{\mathrm{sem}}=\max_{i\in\mathcal{V}_{hit}}\mathtt{DA}\left(F_{\mathrm{L}}^{\mathrm{add}},\mathtt{Proj}\left(q_{\mathrm{3D}}^{k},i\right),F_{\mathrm{I}}^{\mathrm{fet}}\right).$ (5)
@@
Here, $i$ denotes the index of the image view, $F_{\mathrm{L}}^{\mathrm{add}}$ is the aggregated query feature for the $k$-th voxel, and $\mathtt{Proj}\left(q_{\mathrm{3D}}^{k},i\right)$ denotes the projection function mapping the 3D voxel center $q_{\mathrm{3D}}^{k}$ to the $i$-th image view. $\mathcal{V}_{hit}$ is the set of views in which the voxel center is visible. The refined semantic-aware features $F_{\mathrm{L}}^{\mathrm{sem}}$ are then used to replace the original features in $F_{\mathrm{L}}$ at their respective occupied voxel locations $O$, yielding the semantic-aware 3D LiDAR voxel features $V_{\mathrm{L}}\in\mathbb{R}^{\mathrm{D\times H\times W\times C}}$, defined as:
$V_{\mathrm{L}}=\tau\left(F_{\mathrm{L}}^{\mathrm{sem}},F_{\mathrm{L}},O\right),$ (6)
where
$O=\left\{\mathcal{O}_{l}\right\}_{l=\mathrm{1}}^{N_{\mathrm{occ}}}.$ (7)
@@
In this context, $O$ represents the set of 3D locations corresponding to $F_{\mathrm{L}}^{\mathrm{sem}}$, with $N_{\mathrm{occ}}$ occupied voxels in total. Each $\mathcal{O}_{l}$ denotes the 3D index of the $l$-th occupied voxel, and $\tau(\cdot)$ represents the update operation that replaces the original features in $F_{\mathrm{L}}$ with the refined features from $F_{\mathrm{L}}^{\mathrm{sem}}$ at the specified voxel locations.
@@ III-B Late Stage Voxel Fusion
The HCCVF and AF modules are designed for late-stage voxel fusion to refine the performance of 3D semantic occupancy perception.
@@ III-B1 High Classification Confidence Voxel Fusion Module
The HCCVF module, as illustrated in Fig. 2, first employs a dual-branch prediction head to estimate classification confidence scores for both modalities. Top-$K$ ranked voxels from each modality are then selected, denoted as $V_{\mathrm{L}}^{\mathrm{top}}$ and $V_{\mathrm{I}}^{\mathrm{top}}$ for the LiDAR and camera modalities, respectively. Beyond filtering high-confidence occupancy predictions, this mechanism also identifies voxels with conflicting classification confidence scores across modalities, which may indicate semantic ambiguity. These ambiguous voxels are prioritized for refinement and passed through modality-specific learnable projectors: $\mathrm{f_{L}(\cdot)}$ for $V_{\mathrm{L}}^{\mathrm{top}}$ and $\mathrm{f_{C}(\cdot)}$ for $V_{\mathrm{I}}^{\mathrm{top}}$. The projected features are then concatenated along the sequence dimension to form the set $V_{\mathrm{LC}}^{\mathrm{top}}$, defined as:
@@
This concatenated set $V_{\mathrm{LC}}^{\mathrm{top}}$ is subsequently passed through a self-attention module followed by a feed-forward network to obtain the final fused instance features $V_{\mathrm{F}}^{\mathrm{top}}$. Within the self-attention module, learned positional embeddings are incorporated by encoding the XYZ coordinates of the occupancy voxels using a Multi-Layer Perceptron (MLP) and adding these embeddings to the instance features.
@@ III-B2 Adaptive Fusion Module
As depicted in Fig. 2, the AF module employs an adaptive fusion mechanism based on a lightweight 3D convolutional network to enhance the geometric and semantic features at corresponding voxel locations [10]. This results in the fused 3D occupancy voxel features, denoted as $V_{\mathrm{F}}^{\mathrm{ada}}$. The fusion mechanism is formulated as follows:
@@
Here, $\mathtt{Conv}(\cdot)$ denotes the weight encoding network, which primarily consists of a 3D convolutional network and normalization layers. The weight $W_{\mathrm{vis}}$ is computed via a sigmoid function and is used to balance the contribution of semantic features from the camera and geometric features from LiDAR, with values ranging from 0 to 1. Subsequently, the fused instance features $V_{\mathrm{F}}^{\mathrm{top}}$ obtained from the HCCVF module are used to replace part of the features in $V_{\mathrm{F}}^{\mathrm{ada}}$ at the corresponding 3D voxel locations $P$, resulting in the final fused 3D occupancy voxel features $V_{\mathrm{F}}$, defined as:
@@
where
$P=\left\{\mathcal{P}_{h}\right\}_{h=\mathrm{1}}^{N_{\mathrm{LC}}}.$ (12)
In this formulation, $P$ denotes the set of 3D locations corresponding to $V_{\mathrm{F}}^{\mathrm{top}}$, with $N_{\mathrm{LC}}$ being the number of updated voxels. Each $\mathcal{P}_{h}$ represents the $h$-th voxel location in the set, and $\chi(\cdot)$ denotes the replacement operation that updates $V_{\mathrm{F}}^{\mathrm{ada}}$ with $V_{\mathrm{F}}^{\mathrm{top}}$ at the specified locations. Finally, a 3D convolutional decoder is applied to the fused features $V_{\mathrm{F}}$ to produce the final occupancy prediction results.
@@ III-C Loss Function
To train the proposed MS-Occ, the cross-entropy loss $\mathcal{L}_{\text{ce}}$ and the Lovász-Softmax loss $\mathcal{L}_{\text{ls}}$ [32] are utilized to optimize the network. Following [1], affinity losses $\mathcal{L}_{\text{scal}}^{\text{geo}}$ and $\mathcal{L}_{\text{scal}}^{\text{sem}}$ are also employed to optimize scene-level and class-level metrics, respectively. In addition, explicit depth supervision $\mathcal{L}_{\text{d}}$ [33] is introduced to guide the training of the depth-aware view transformation module. An auxiliary mask supervision is applied using focal loss, denoted as $\mathcal{L}_{\text{mask}}$. The total mask loss is formulated as:
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\mathrm{ls}}+\mathcal{L}_{\mathrm{scal}}^{\mathrm{geo}}+\mathcal{L}_{\mathrm{scal}}^{\mathrm{sem}}+\mathcal{L}_{\mathrm{d}}+\mathcal{L}_{\mathrm{mask}}.$ (13)
where
@@
Here, $\hat{V_{\mathrm{L}}}$ and $\hat{V_{\mathrm{I}}}$ denote the classification predictions for the LiDAR and camera modalities, respectively, while $\mathcal{V_{\mathrm{L}}}$ and $\mathcal{V_{\mathrm{I}}}$ represent their corresponding ground truth voxel labels.
TABLE I: Comparison between the proposed MS-Occ and state-of-the-art methods on the 3D semantic occupancy prediction task using the nuScenes-OpenOccupancy validation set [34, 10]. “C” and “L” denote camera and LiDAR modalities, respectively.
@@
MonoScene (CVPR 2022) [1] C 24.0 7.3 4.0 0.4 8.0 8.0 2.9 0.3 1.2 0.7 4.0 4.4 27.7 5.2 15.1 11.3 9.0 14.9
TPVFormer (CVPR 2023) [18] C 15.3 7.8 9.3 4.1 11.3 10.1 5.2 4.3 5.9 5.3 6.8 6.5 13.6 9.0 8.3 8.0 9.2 8.2
BEVFormer (ECCV 2022) [3] C 30.5 16.7 14.2 6.5 23.4 28.2 8.6 10.7 6.4 4.0 11.2 17.7 37.2 18.0 22.8 22.1 13.8 22.2
SurroundOcc (ICCV 2023) [2] C 31.4 20.3 20.5 11.6 28.1 30.8 10.7 15.1 14.0 12.0 14.3 22.2 37.2 23.7 24.4 22.7 14.8 21.8
OccFormer (ICCV 2023) [21] C 29.9 20.1 21.1 11.3 28.2 30.3 10.6 15.7 14.4 11.2 14.0 22.6 37.3 22.4 24.9 23.5 15.2 21.1
LMSCNet (3DV 2020) [35] L 27.3 11.5 12.4 4.2 12.8 12.1 6.2 4.7 6.2 6.3 8.8 7.2 24.2 12.3 16.6 14.1 13.9 22.2
L-CONet (ICCV 2023) [2] L 30.9 15.8 17.5 5.2 13.3 18.1 7.8 5.4 9.6 5.6 13.2 13.6 34.9 21.5 22.4 21.7 19.2 23.5
M-CONet (CVPR 2023) [10] C&L 29.5 20.1 23.3 13.3 21.2 24.3 15.3 15.9 18.0 13.3 15.3 20.7 33.2 21.0 22.5 21.5 19.6 23.2
Co-Occ (IEEE RA-L 2024) [11] C&L 30.6 21.9 26.5 16.8 22.3 27.0 10.1 20.9 20.7 14.5 16.4 21.6 36.9 23.5 25.5 23.7 20.5 23.5
OccGen (ECCV 2024) [12] C&L 30.3 22.0 24.9 16.4 22.5 26.1 14.0 20.1 21.6 14.6 17.4 21.9 35.8 24.5 24.7 24.0 20.5 23.5
OccLoff (WACV 2025) [9] C&L 31.4 22.9 26.7 17.2 22.6 26.9 16.4 22.6 24.7 16.4 16.3 22.0 37.5 22.3 25.3 23.9 21.4 24.2
ML-Occ(ours) C&L 32.1 25.3 29.4 23.8 24.2 28.7 17.5 27.5 29.0 21.1 16.9 25.3 33.4 23.5 24.4 24.6 27.6 28.4
IV Experiments
IV-A Dataset and Metrics
MS-Occ is evaluated on the nuScenes-OpenOccupancy dataset [34, 10], a widely accepted large-scale benchmark for 3D semantic occupancy prediction in autonomous driving. The dataset comprises 1,000 driving scenarios, consisting of 700 training scenes, 150 validation scenes, and 150 testing scenes. Each sample is annotated with 3D occupancy ground truth provided by [10], using a voxel resolution of (0.2 m, 0.2 m, 0.2 m), forming a dense voxel grid of size $(512,512,40)$. A total of 17 classes are defined, including 16 semantic categories and one class representing free space.
Following [10], scene completion intersection over union (IoU) is used to evaluate geometric accuracy, while mean IoU (mIoU) is adopted to assess semantic accuracy. These metrics are defined as:
$IoU=\frac{TP}{TP+FP+FN},$ (15)
and
$mIoU=\frac{1}{S}\sum_{i=1}^{S}\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}},$ (16)
where $TP$, $FP$, and $FN$ denote the number of true positives, false positives, and false negatives, respectively, and $S$ is the number of semantic classes.
IV-B Implementation Details
For the camera branch, input images are resized to a resolution of $1600\times 900$ and processed using a pretrained ResNet-50 [36] backbone with an FPN [37] for 2D feature extraction. For the LiDAR branch, 10 LiDAR sweeps are voxelized and encoded using a voxel-based encoder for spatial feature learning [38, 39]. The Gaussian-Geo module uses a Gaussian kernel configuration of size 6, and the enhanced depth heatmaps are processed using $1\times 1$ convolutional layers. The fused depth features are subsequently encoded by a ResNet-50 [36] backbone. In the late fusion stage, a 3D decoder with an upsampling stride of 2 is applied to recover high-resolution volumetric features.
The model is implemented using the MMDetection3D framework [40] and trained for 24 epochs with a batch size of 8 distributed across 8 NVIDIA A100 GPUs. During training, the AdamW optimizer [41] is employed with a weight decay of 0.01 and an initial learning rate of $3\times 10^{-4}$.

### v2

@@ III-A2 Semantic-aware Module
As shown in Fig. 4, given the voxel features $F_{\mathrm{L}}$, only a limited number of voxels are actually occupied by LiDAR point clouds. For each occupied voxel at location $(x,y,z)$, where $x\in\left\{0,1,\cdots,W-1\right\}$, $y\in\left\{0,1,\cdots,H-1\right\}$, and $z\in\left\{0,1,\cdots,D-1\right\}$, the voxel center is projected onto the depth-aware image features of each view, denoted as $F_{\mathrm{I}}^{\mathrm{geo}}$. The corresponding 2D image features $F_{\mathrm{I}}^{\mathrm{fet}}$ are then extracted at these projected points and combined with the original occupied voxel features $F_{\mathrm{L}}^{\mathrm{occ}}$ through element-wise addition. The resulting aggregated features $F_{\mathrm{L}}^{\mathrm{add}}$ serve as queries and interact with the multi-scale image features $F_{\mathrm{I}}^{\mathrm{fet}}$ of the respective image views via a deformable attention module [DeformableDETR, bevformer]. This process is performed independently for each image view:
@@
$\mathtt{DA}\left(u,r,e\right)=\sum_{a=1}^{N_{\mathrm{head}}}W_{a}\sum_{b=1}^{N_{\mathrm{key}}}A_{ab}\cdot W_{a}^{\prime}e\left(r+\Delta r_{ab}\right).$ (4)
@@
Here, $u$ and $e$ denote the queries and keys, respectively, $W_{a}$ and $W_{a}^{\prime}$ are learnable projection matrices, $A_{ab}$ represents the attention weights, $r$ is the 2D reference point onto which the voxel center is projected, and $\Delta r_{ab}$ denotes the positional offset. Due to the overlapping fields of multiple views, a voxel may be projected onto several image views. To consolidate the updated features across views, max-pooling is applied, resulting in the output feature $F_{\mathrm{L}}^{\mathrm{sem}}=\left\{F_{\mathrm{L},k}^{\mathrm{sem}}\right\}_{k=1}^{N_{\mathrm{v}}}$, where $N_{\mathrm{v}}$ is the number of occupied voxels. Each updated voxel feature $F_{\mathrm{L},k}^{\mathrm{sem}}$ is defined as:
$F_{\mathrm{L},k}^{\mathrm{sem}}=\max_{j\in\mathcal{V}_{hit}}\mathtt{DA}\left(F_{\mathrm{L}}^{\mathrm{add}},\mathtt{Proj}\left(q_{\mathrm{3D}}^{k},j\right),F_{\mathrm{I}}^{\mathrm{fet}}\right).$ (5)
@@
Here, $j$ denotes the index of the image view, $F_{\mathrm{L}}^{\mathrm{add}}$ is the aggregated query feature for the $k$-th voxel, and $\mathtt{Proj}\left(q_{\mathrm{3D}}^{k},j\right)$ denotes the projection function mapping the 3D voxel center $q_{\mathrm{3D}}^{k}$ to the $j$-th image view. $\mathcal{V}_{hit}$ is the set of views in which the voxel center is visible. The refined semantic-aware features $F_{\mathrm{L}}^{\mathrm{sem}}$ are then used to replace the original features in $F_{\mathrm{L}}$ at their respective occupied voxel locations $O$, yielding the semantic-aware 3D LiDAR voxel features $V_{\mathrm{L}}$, defined as:
$V_{\mathrm{L}}$ $=\tau\left(F_{\mathrm{L}}^{\mathrm{sem}},F_{\mathrm{L}},O\right)\in\mathbb{R}^{\mathrm{D\times H\times W\times C}},$ (6)
$O$ $=\left\{\mathcal{O}_{l}\right\}_{l=1}^{N_{\mathrm{occ}}}.$ (7)
@@
In this context, $O$ represents the set of 3D locations corresponding to $F_{\mathrm{L}}^{\mathrm{sem}}$, with $N_{\mathrm{occ}}$ occupied voxels in total. Each $\mathcal{O}_{l}$ denotes the 3D index of the $l$-th occupied voxel, and $\tau(\cdot)$ represents the update operation.
@@ III-B Late Stage Voxel Fusion
The HCCVF and AF modules are designed to perform late-stage voxel fusion to refine 3D semantic occupancy perception.
@@ III-B1 High Classification Confidence Voxel Fusion Module
The HCCVF module, as illustrated in Fig. 2, first employs a dual-branch prediction head to estimate classification confidence scores for both modalities. Top-$K$ ranked voxels from each modality are then selected, denoted as $V_{\mathrm{L}}^{\mathrm{top}}$ and $V_{\mathrm{I}}^{\mathrm{top}}$ for the LiDAR and camera modalities, respectively. Beyond filtering high-confidence occupancy predictions, this mechanism also identifies voxels with conflicting classification confidence scores across modalities, which may indicate semantic ambiguity. These ambiguous voxels are prioritized for refinement and passed through modality-specific learnable projectors: $\mathrm{f_{L}(\cdot)}$ for $V_{\mathrm{L}}^{\mathrm{top}}$ and $\mathrm{f_{C}(\cdot)}$ for $V_{\mathrm{I}}^{\mathrm{top}}$. The projected features are then concatenated along the sequence dimension to form the set $V_{\mathrm{LC}}^{\mathrm{top}}$:
@@
This concatenated set is subsequently passed through a self-attention module followed by a feed-forward network to obtain the final fused instance features $V_{\mathrm{F}}^{\mathrm{top}}$. Within the self-attention module, learned positional embeddings are incorporated by encoding the XYZ coordinates of the occupancy voxels using a Multi-Layer Perceptron (MLP) and adding them to the instance features.
@@ III-B2 Adaptive Fusion Module
As depicted in Fig. 2, the AF module employs an adaptive fusion mechanism based on a lightweight 3D convolutional network to enhance the geometric and semantic features at corresponding voxel locations [wang2023openoccupancy]. This results in the fused 3D occupancy voxel features, denoted as $V_{\mathrm{F}}^{\mathrm{ada}}$. The fusion mechanism is formulated as:
@@
Here, $\mathtt{Conv}(\cdot)$ denotes the weight encoding network, which primarily consists of a 3D convolutional network and normalization layers. The weight $W_{\mathrm{vis}}$ is computed via a sigmoid function and is used to balance the contribution of semantic features from the camera and geometric features from LiDAR. Subsequently, the fused instance features $V_{\mathrm{L}}^{\mathrm{top}}$ from the HCCVF module replace part of the features in $V_{\mathrm{F}}^{\mathrm{ada}}$ at the corresponding 3D voxel locations $P$, resulting in the final fused 3D occupancy voxel features $V_{\mathrm{F}}$:
@@
$P$ $=\left\{\mathcal{P}_{h}\right\}_{h=1}^{N_{\mathrm{LC}}}.$ (12)
In this formulation, $P$ denotes the set of 3D locations corresponding to $V_{\mathrm{F}}^{\mathrm{top}}$, and $\chi(\cdot)$ denotes the replacement operation. Finally, a 3D convolutional decoder is applied to the fused features $V_{\mathrm{F}}$ to produce the final occupancy prediction results.
@@ III-C Loss Function
To train the proposed MS-Occ, the cross-entropy loss $\mathcal{L}_{\text{ce}}$ and the Lovász-Softmax loss $\mathcal{L}_{\text{ls}}$ [lovasz-softmax] are utilized. Following [monoscene], affinity losses $\mathcal{L}_{\text{scal}}^{\text{geo}}$ and $\mathcal{L}_{\text{scal}}^{\text{sem}}$ are also employed. In addition, explicit depth supervision $\mathcal{L}_{\text{d}}$ [bevdepth] is introduced to guide the depth-aware view transformation module. An auxiliary mask supervision is applied using focal loss, denoted as $\mathcal{L}_{\text{mask}}$. The total loss is formulated as:
$\mathcal{L}_{\mathrm{total}}$ $=\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\mathrm{ls}}+\mathcal{L}_{\mathrm{scal}}^{\mathrm{geo}}+\mathcal{L}_{\mathrm{scal}}^{\mathrm{sem}}+\mathcal{L}_{\mathrm{d}}+\mathcal{L}_{\mathrm{mask}},$ (13)
@@
where $\hat{V_{\mathrm{L}}}$ and $\hat{V_{\mathrm{I}}}$ are the classification predictions, while $\mathcal{V_{\mathrm{L}}}$ and $\mathcal{V_{\mathrm{I}}}$ represent their corresponding ground truth voxel labels.
IV Experiments
IV-A Dataset and Metrics
Our method is evaluated on two widely accepted, large-scale benchmarks for 3D semantic occupancy prediction. The primary dataset is nuScenes-OpenOccupancy [wang2023openoccupancy, nuscenes2019], which comprises 1,000 driving scenarios, annotated with 3D occupancy ground truth provided by [wang2023openoccupancy], using a dense voxel grid of size $(512,512,40)$. A total of 17 classes are defined, including 16 semantic categories and one class representing free space. To assess generalization capability, we further evaluate on the SemanticKITTI dataset [38], comprising 22 sequences with a voxel grid of size $(256,256,32)$, annotated with 21 total classes (19 semantic, 1 free, and 1 unknown). SemanticKITTI [38] presents a more challenging benchmark featuring denser, complex, and unstructured natural environments with significant terrain and vegetation variations. Both benchmarks utilize a voxel resolution of (0.2 m, 0.2 m, 0.2 m). Following [wang2023openoccupancy], scene completion intersection over union (IoU) is used to evaluate geometric accuracy, while mean IoU (mIoU) is adopted for semantic accuracy.
IV-B Implementation Details
To ensure a fair comparison, our implementation and key hyperparameter choices closely follow the experimental settings established by prior benchmarks [wang2023openoccupancy, Co-Occ]. For the camera branch, input images are resized to $1600\times 900$ and processed using a pretrained ResNet-50 [Resnet] backbone with an FPN [lin2017fpn]. For the LiDAR branch, 10 LiDAR sweeps are voxelized and encoded using a voxel-based encoder [Zhou_Tuzel_VoxelNet, SECOND]. To balance predictive accuracy and computational efficiency, the Gaussian kernel size $\sigma$ is set as 6, a choice validated through ablation studies (Table IV). The enhanced depth heatmaps from the Gaussian-Geo module are processed using $1\times 1$ convolutional layers, and the fused depth features are subsequently encoded by a ResNet-50 [Resnet] backbone. In the late fusion stage, a 3D decoder with an upsampling stride of 2 is applied. The model is implemented using MMDetection3D [contributors2020mmdetection3d] and trained for 24 epochs with a batch size of 8 on 8 NVIDIA A100 GPUs. We use the AdamW optimizer [adamw] with a weight decay of 0.01 and an initial learning rate of $3\times 10^{-4}$.
TABLE I: Comparison with state-of-the-arts on nuScenes-OpenOccupancy validation set [nuscenes2019, wang2023openoccupancy]. “C” and “L” denote camera and LiDAR.
@@ IV-B Implementation Details
TPVFormer (CVPR 2023) [TPVFormer] C 15.3 7.8 5.9 4.1 4.3 5.3 9.3 9.2 10.1 11.3 6.5 6.8 5.2 13.6 8.3 8.0 9.0 8.2
SurroundOcc (ICCV 2023) [surroundocc] C 31.4 20.3 14.0 11.6 15.1 12.0 20.5 14.8 30.8 28.1 22.2 14.3 10.7 37.2 24.4 22.7 23.7 21.8
OccFormer (ICCV 2023) [zhang2023occformer] C 29.9 20.1 14.4 11.3 15.7 11.2 21.1 15.2 30.3 28.2 22.6 14.0 10.6 37.3 24.9 23.5 22.4 21.1
LMSCNet (3DV 2020) [lmscnet] L 27.3 11.5 6.2 4.2 4.7 6.3 12.4 13.9 12.1 12.8 7.2 8.8 6.2 24.2 16.6 14.1 12.3 22.2
L-CONet (ICCV 2023) [wang2023openoccupancy] L 30.9 15.8 9.6 5.2 5.4 5.6 17.5 19.2 18.1 13.3 13.6 13.2 7.8 34.9 22.4 21.7 21.5 23.5
M-CONet (ICCV 2023) [wang2023openoccupancy] C&L 29.5 20.1 18.0 13.3 15.9 13.3 23.3 19.6 24.3 21.2 20.7 15.3 15.3 33.2 22.5 21.5 21.0 23.2
Co-Occ (IEEE RA-L 2024) [Co-Occ] C&L 30.6 21.9 20.7 16.8 20.9 14.5 26.5 20.5 27.0 22.3 21.6 16.4 10.1 36.9 25.5 23.7 23.5 23.5
OccGen (ECCV 2024) [OccGen] C&L 30.3 22.0 21.6 16.4 20.1 14.6 24.9 20.5 26.1 22.5 21.9 17.4 14.0 35.8 24.7 24.0 24.5 23.5
OccLoff (WACV 2025) [occloff] C&L 31.4 22.9 24.7 17.2 22.6 16.4 26.7 21.4 26.9 22.6 22.0 16.3 16.4 37.5 25.3 23.9 22.3 24.2
MS-Occ(ours) C&L 32.1 25.3 29.0 23.8 27.5 21.1 29.4 27.6 28.7 24.2 25.3 16.9 17.5 33.4 24.4 24.6 23.5 28.4
TABLE II: Comparison with state-of-the-arts on the SemanticKITTI validation set [semantickitti]. Classes are grouped to highlight performance on VRUs, Small Objects, and Complex Structures. “C” and “L” denote camera and LiDAR.
Methods Modality mIoU
$\blacksquare$ person
$\blacksquare$ bicyclist
$\blacksquare$ motorcyclist.
$\blacksquare$ bicycle
$\blacksquare$ motorcycle
$\blacksquare$ traf.-sign
$\blacksquare$ car
$\blacksquare$ truck
$\blacksquare$ other-veh.
$\blacksquare$ vegetation
$\blacksquare$ trunk
$\blacksquare$ fence
$\blacksquare$ road
$\blacksquare$ sidewalk
$\blacksquare$ parking
$\blacksquare$ other-grnd
$\blacksquare$ building
$\blacksquare$ terrain
$\blacksquare$ pole
JS3C-Net (AAAI 2021) [JS3C-Net] C 10.31 0.67 0.27 0.00 0.00 0.00 1.45 24.65 4.41 6.15 18.11 4.33 3.94 50.49 23.74 11.94 0.07 15.03 26.86 3.77
MonoScene (CVPR 2022) [monoscene] C 11.08 1.86 1.20 0.00 0.61 0.45 2.25 23.26 6.98 1.48 17.89 2.81 5.84 56.52 26.72 14.27 0.46 14.09 29.64 4.14
VoxFormer (CVPR 2023) [li2023voxformer] C 12.35 1.78 3.32 0.00 0.59 0.51 4.18 25.79 5.63 3.77 24.39 5.08 7.64 54.76 26.35 15.50 0.70 17.65 29.96 7.11
LMSCNet (3DV 2020) [lmscnet] L 6.70 0.00 0.00 0.00 0.0 0.0 0.00 18.33 0.00 0.0 13.66 0.02 1.21 40.68 18.22 4.38 0.00 10.31 20.54 0.00
M-CONet (CVPR 2023) [wang2023openoccupancy] C&L 17.25 2.07 2.38 0.00 0.66 2.2 14.09 33.25 19.46 5.99 32.56 16.01 12.25 58.00 29.44 19.33 0.02 27.85 37.31 14.88
OccFusion (IEEE T-IV 2024) [occfusion] C&L 21.64 2.33 3.07 0.00 2.00 3.39 14.36 45.72 19.41 7.92 41.21 19.18 15.74 65.73 36.46 21.35 0.00 39.32 46.28 27.60
OccLoff (WACV 2025) [occloff] C&L 22.62 3.88 4.35 0.00 2.08 3.91 16.47 46.44 20.38 8.72 41.20 20.06 15.86 66.25 43.51 21.07 0.57 41.23 46.21 27.60
MS-Occ (Ours) C&L 24.08 5.32 6.59 0.00 3.96 4.65 18.61 47.78 22.82 10.86 46.44 24.7 18.6 64.89 42.65 20.81 0.51 44.67 46.15 27.44
TABLE III: Comparison of accuracy and efficiency with state-of-the-art methods on the nuScenes validation set. All Frames Per Second (FPS) benchmarks were conducted on a single NVIDIA A100 GPU for a fair comparison.
Method   Params (M)   GFLOPs   FPS   IoU   mIoU
M-OpenOccupancy 123.37 3045.11 2.3 29.1 15.1
M-CONet 123.45 3089.47 1.6 29.5 20.1
MS-Occ (Ours) 100.8 3617.9 1.3 32.1 25.3
Figure 5: Visualization results on the nuScenes-OpenOccupancy dataset [nuscenes2019, wang2023openoccupancy]. The leftmost column shows the surround-view images. The next three columns present 3D semantic occupancy predictions from M-CONet, MS-Occ (ours), and the ground truth. Please zoom in for finer details.